# Diffusion-Reward Adversarial Imitation Learning

### v1

@@ Abstract
Imitation learning aims to learn a policy from observing expert demonstrations without access to reward signals from environments. Generative adversarial imitation learning (GAIL) formulates imitation learning as adversarial learning, employing a generator policy learning to imitate expert behaviors and discriminator learning to distinguish the expert demonstrations from agent trajectories. Despite its encouraging results, GAIL training is often brittle and unstable. Inspired by the recent dominance of diffusion models in generative modeling, we propose Diffusion-Reward Adversarial Imitation Learning (DRAIL), which integrates a diffusion model into GAIL, aiming to yield more robust and smoother rewards for policy learning. Specifically, we propose a diffusion discriminative classifier to construct an enhanced discriminator, and design diffusion rewards based on the classifier’s output for policy learning. Extensive experiments are conducted in navigation, manipulation, and locomotion, verifying DRAIL’s effectiveness compared to prior imitation learning methods. Moreover, additional experimental results demonstrate the generalizability and data efficiency of DRAIL. Visualized learned reward functions of GAIL and DRAIL suggest that DRAIL can produce more robust and smoother rewards.
@@ 1 Introduction
We extensively compare our proposed framework DRAIL with behavioral cloning [40], Diffusion Policy [38, 6], and AIL methods, e.g., GAIL [19] , WAIL [2], and DiffAIL[55] , in diverse continuous control domains, including navigation, robot arm manipulation, locomotion, and games. This collection of tasks includes environments with high-dimensional continuous state and action spaces, as well as covers both vectorized and image-based states. The experimental results show that our proposed framework consistently outperforms the prior methods or achieves competitive performance. Moreover, DRAIL exhibits superior performance in generalizing to states or goals unseen from the expert ‘demonstrations. When varying the amounts of available expert data, DRAIL demonstrates the best data efficiency. At last, the visualized learned reward functions show that DRAIL captures more robust and smoother rewards compared to GAIL.
@@ 2 Related work
Adversarial Imitation Learning (AIL). AIL methods aim to directly match the state-action distributions of an agent and an expert through adversarial training. Generative adversarial imitation learning (GAIL) [19] and its extensions [52, 28, 58, 22, 55] train a generator policy to imitate expert behaviors and a discriminator to differentiate between the expert and the generator’s state-action pair distributions, which resembles the idea of generative adversarial networks (GANs) [14]. Thanks to its simplicity and effectiveness, GAIL has been widely applied to various domains [3, 41, 25]. Over the past years, researchers and practitioners have proposed numerous improvements to enhance GAIL’s sample efficiency, scalability, and robustness [36], including modifications to discriminator’s loss function [11], extensions to off-policy RL algorithms [27], and exploration of various similarity measures [11, 2, 8]. In this work, we propose to use the diffusion model as a discriminator in GAIL.
@@ 4.2 Diffusion discriminative classifier
Our goal is to yield a diffusion model reward given an agent state-action pair without going through the entire diffusion generation process. To this end, we propose to leverage the training procedure of DDPM and develop a mechanism to provide learning signals to a policy using a single denoising step. Our key insight is to leverage the derivations developed by Song et al. [48], Kingma et al. [23], which suggest that the diffusion loss, i.e., the difference between the predicted noise and the injected noise, indicates how well the data fits the target distribution since the diffusion loss is the upper bound of the negative log-likelihood of data in the target distribution. In this work, we propose calculating “realness” rewards based on the diffusion loss computed by denoising the state-action pairs from the policy, which indicates how well the state-action pairs fit the expert behavior distributions. We formulate the diffusion loss $\mathcal{L}_{\text{diff}}$ as follows:
@@
where $\mathbf{c}\in\{\mathbf{c}^{+},\mathbf{c}^{-}\}$, and the real label $\mathbf{c}^{+}$ corresponds to the condition for fitting expert data while the fake label $\mathbf{c}^{-}$ corresponds to agent data. We implement $\mathbf{c}^{+}$ as $\mathbf{1}$ and $\mathbf{c}^{-}$ as $\mathbf{0}$. Subsequently, given a state-action pair $(\mathbf{s},\mathbf{a})$, $\mathcal{L}_{\text{diff}}(\mathbf{s},\mathbf{a},\mathbf{c}^{+})$ measures how well $(\mathbf{s},\mathbf{a})$ fits the expert distribution and $\mathcal{L}_{\text{diff}}(\mathbf{s},\mathbf{a},\mathbf{c}^{-})$ measures how well $(\mathbf{s},\mathbf{a})$ fits the agent distribution [^1] For simplicity, we will use the notations $\mathcal{L}^{+}_{\text{diff}}$ and $\mathcal{L}^{-}_{\text{diff}}$ to represent $\mathcal{L}_{\text{diff}}(\mathbf{s},\mathbf{a},\mathbf{c}^{+})$ and $\mathcal{L}_{\text{diff}}(\mathbf{s},\mathbf{a},\mathbf{c}^{-})$, respectively, in the rest of the paper. Additionally, $\mathcal{L}^{+,-}_{\text{diff}}$ denotes $\mathcal{L}^{+}_{\text{diff}}$ and $\mathcal{L}^{-}_{\text{diff}}$ given a state-action pair.. That said, given state-action pairs sampled from expert demonstration, $\mathcal{L}^{+}_{\text{diff}}$ should be close to 00, and $\mathcal{L}^{-}_{\text{diff}}$ should be a large value; on the contrary, given agent state-action pairs, $\mathcal{L}^{+}_{\text{diff}}$ should be a large value and $\mathcal{L}^{-}_{\text{diff}}$ should close to 00.
@@ 5.2 Baselines
  • Diffusion Adversarial Imitation Learning (DiffAIL) [54] integrates a diffusion model into AIL by using the diffusion model loss to provide reward $e^{-\mathcal{L}_{\text{diff}}}$; however, this does not resemble the binary classification formulation in GAIL. In contrast, our discriminator is trained to perform binary classification as in GAIL, allowing for providing better learning signal especially when the agent policy gets closer to the expert, detailed in Section A.
@@ 5.3 Experimental results
We present the success rates (Maze, FetchPush, HandRotate, AntReach) and return (Walker, CarRacing) of all the methods with regards to environment steps in Figure 4. Each task was trained using five different random seeds. Note that BC and Diffusion Policy are offline imitation learning algorithms, meaning they cannot interact with the environment, so their performances are represented as horizontal lines. Detailed information on model architectures, training, and evaluation can be found in Section E and Section F.
@@ 5.4 Generalizability
The results of FetchPush under 1×1$\times$1 ×, 1.25×1.25$\times$1.25 ×, 1.75×1.75$\times$1.75 ×, and 2.0×2.0$\times$2.0 × noise level are presented in Figure 5. Across all noise levels, approaches utilizing the diffusion model generally exhibit better performance. Notably, our proposed DRAIL demonstrates the highest robustness towards noisy environments. Even at the highest noise level of 2.00×2.$0$, DRAIL maintains a success rate of over $95\%$, surpassing the best-performing baseline, Diffusion Policy, which achieves only around a $79.20\%$ success rate. In contrast, DiffAIL experiences failures in $2$ out of the $5$ seeds, resulting in a high standard deviation (mean: $40.90$, standard deviation: $47.59$), despite thorough exploration of various configuration settings.
The extended generalization experiments results in Maze, FetchPush, HandRotate, AntReach, and the new task FetchPick are presented in Section C. Overall, our method outperforms or performs competitively against the best-performing baseline, demonstrating its superior generalization ability.
@@ 5.5 Data efficiency
Figure 7: Reward function visualization. We present visualizations of the learned reward values by the discriminative classifier of GAIL and the diffusion discriminative classifier of our DRAIL. The target expert demonstration for imitation is depicted in (7(a)), which is a discontinuous sine function. The reward distributions of GAIL and our DRAIL are illustrated in (7(b)) and (7(c)), respectively.
@@ 6 Conclusion
This work proposes a novel adversarial imitation learning framework that integrates a diffusion model into generative adversarial imitation learning (GAIL). Specifically, we propose a diffusion discriminative classifier that employs a diffusion model to construct an enhanced discriminator, yielding more robust and smoother rewards. Then, we design diffusion rewards based on the classifier’s output for policy learning. Extensive experiments in navigation, manipulation, locomotion, and game justify our proposed framework’s effectiveness, generalizability, and data efficiency.
@@ Appendix A Relation to DiffAIL
Experimental setup. Wang et al. [54] evaluated DiffAIL and prior methods in locomotion tasks; in contrast, our work extensively compares our proposed framework DRAIL with various existing methods in various domains, including navigation (Maze and AntReach), locomotion (Walker and AntReach), robot arm manipulation (FetchPush and FetchPick), robot arm dexterous (HandRotate), and games (CarRacing). Additionally, we present experimental results on generalization to unseen states and goals on the goal-oriented tasks,
@@ C.1 Experiment settings
To show our approach’s better generalization capabilities, we extend the environment scenarios following the setting stated in Lee et al. [30]: (1) In Maze main experiment, the initial and goal states of the expert dataset only constitute $50\%$ of the potential initial and goal states. In the generalization experiment, we gather expert demonstrations from some lower and higher coverage: $25\%$, $75\%$, and $100\%$. (2) In FetchPick, FetchPush, and HandRotate main experiments, the demonstrations are collected in a lower noise level setting, 1×1$\times$1 ×. Yet, the agent is trained within an environment incorporating 1.5×1.5$\times$1.5 × noise, which is 1.5 times larger noise than the collected expert demonstration, applied to the starting and target block positions. In the generalization experiment, we train agents in different noise levels: 1×1$\times$1 ×, 1.25×1.25$\times$1.25 ×, 1.75×1.75$\times$1.75 ×, 2.0×2.0$\times$2.0 ×. (3) In AntReach main experiment, $0.03$ random noise is added to the initial pose during policy learning. In AntReach generalization experiment, we train agents in different noise levels: 00, $0.01$, $0.05$.
@@ C.2 Experiment results
(q) AntReach 0.00
(r) AntReach 0.01
(s) AntReach 0.03
(t) AntReach 0.05
Figure 8: Extended results of generalization experiments. Maze is evaluated with different coverages of state and goal locations in the expert demonstrations, while FetchPick, FetchPush, HandRotate, and AntReach environments are evaluated in environments of different noise levels. The number indicates the amount of additional noise in agent learning compared to that in the expert demonstrations, with more noise requiring harder generalization. The noise level rises from left to right.
@@
AntReach. Under a $0.03$ noise level, the overall success rate of our method, and the baselines drop dramatically, as illustrated in Figure 4(d). Despite this, our DRAIL maintains a higher success rate of $39.02\%$, surpassing other baselines that achieve success rates lower than $35\%$.
Appendix D Converged performance

### v4

@@ Abstract
Imitation learning aims to learn a policy from observing expert demonstrations without access to reward signals from environments. Generative adversarial imitation learning (GAIL) formulates imitation learning as adversarial learning, employing a generator policy learning to imitate expert behaviors and discriminator learning to distinguish the expert demonstrations from agent trajectories. Despite its encouraging results, GAIL training is often brittle and unstable. Inspired by the recent dominance of diffusion models in generative modeling, we propose Diffusion-Reward Adversarial Imitation Learning (DRAIL), which integrates a diffusion model into GAIL, aiming to yield more robust and smoother rewards for policy learning. Specifically, we propose a diffusion discriminative classifier to construct an enhanced discriminator, and design diffusion rewards based on the classifier’s output for policy learning. Extensive experiments are conducted in navigation, manipulation, and locomotion, verifying DRAIL’s effectiveness compared to prior imitation learning methods. Moreover, additional experimental results demonstrate the generalizability and data efficiency of DRAIL. Visualized learned reward functions of GAIL and DRAIL suggest that DRAIL can produce more robust and smoother rewards. Project page: https://nturobotlearninglab.github.io/DRAIL/
@@ 1 Introduction
We extensively compare our proposed framework DRAIL with behavioral cloning [44], Diffusion Policy [42, 6], and AIL methods, e.g., GAIL [21], WAIL [2], and DiffAIL[59], in diverse continuous control domains, including navigation, robot arm manipulation, locomotion, and games. This collection of tasks includes environments with high-dimensional continuous state and action spaces, as well as covers both vectorized and image-based states. The experimental results show that our proposed framework consistently outperforms the prior methods or achieves competitive performance. Moreover, DRAIL exhibits superior performance in generalizing to states or goals unseen from the expert’s demonstrations. When varying the amounts of available expert data, DRAIL demonstrates the best data efficiency. At last, the visualized learned reward functions show that DRAIL captures more robust and smoother rewards compared to GAIL.
@@ 2 Related work
Adversarial Imitation Learning (AIL). AIL methods aim to directly match the state-action distributions of an agent and an expert through adversarial training. Generative adversarial imitation learning (GAIL) [21] and its extensions [56, 31, 62, 25, 59, 23] train a generator policy to imitate expert behaviors and a discriminator to differentiate between the expert and the generator’s state-action pair distributions, which resembles the idea of generative adversarial networks (GANs) [16]. Thanks to its simplicity and effectiveness, GAIL has been widely applied to various domains [3, 45, 28]. Over the past years, researchers have proposed numerous improvements to enhance GAIL’s sample efficiency, scalability, and robustness [40], including modifications to discriminator’s loss function [12], extensions to off-policy RL algorithms [30], addressing reward bias [31], and exploration of various similarity measures [12, 2, 8, 11]. Another line of work avoids adversarial training, such as IQ-Learn [15], which learns a Q-function that implicitly represents the reward function and policy. In this work, we propose to use the diffusion model as a discriminator in GAIL.
Diffusion Model-Based Approaches in Reinforcement Learning. Diffuser [57] and Nuti et al. [39] apply diffusion models to reinforcement learning (RL) and reward learning, their settings differ significantly from ours. Diffuser [57] is a model-based RL method that requires trajectory-level reward information, which differs from our setting, i.e., imitation learning, where obtaining rewards is not possible. Nuti et al. [39] focus on learning a reward function, unlike imitation learning, whose goal is to obtain a policy. Hence, Nuti et al. [39] neither present policy learning results in the main paper nor compare their method to imitation learning methods. Moreover, they focus on learning from a fixed suboptimal dataset, AIL approaches and our method are designed to learn from agent data that continually change as the agents learn.
@@ 3.2 Diffusion models
Beyond generative tasks, diffusion models have also been successfully applied in other areas, including image classification and imitation learning. Diffusion Classifier [35] demonstrates that conditional diffusion models can estimate class-conditional densities for zero-shot classification. In imitation learning, diffusion models have been used to improve Behavioral Cloning (DBC) [5] by using diffusion model to model expert state-action pairs. Similarly, DiffAIL [58] extends GAIL [21] by employing diffusion models to represent the expert’s behavior and incorporating the diffusion loss into the discriminator’s learning process. However, DiffAIL’s use of an unconditional diffusion model limits its ability to distinguish between expert and agent state-action pairs. We provide a detailed explanation of its limitation in Section 4.3 and Section A.
@@ 4.2 Diffusion discriminative classifier
Our goal is to yield a diffusion model reward given an agent state-action pair without going through the entire diffusion generation process. Inspired by previous work [5, 59], we extract the learning signal from a portion of the diffusion denoising steps, rather than using the entire process. Building on these insights, we adapt the training procedure of DDPM to develop a mechanism that provides a binary classification signal using just one denoising step.
Our key insight is to leverage the derivations developed by Song et al. [52], Kingma et al. [26], which suggest that the diffusion loss, i.e., the difference between the predicted noise and the injected noise, indicates how well the data fits the target distribution since the diffusion loss is the upper bound of the negative log-likelihood of data in the target distribution. In this work, we propose calculating “realness” rewards based on the diffusion loss computed by denoising the state-action pairs from the policy, which indicates how well the state-action pairs fit the expert behavior distributions. We formulate the diffusion loss $\mathcal{L}_{\text{diff}}$ as follows:
@@
where $\mathbf{c}\in\{\mathbf{c}^{+},\mathbf{c}^{-}\}$, and the real label $\mathbf{c}^{+}$ corresponds to the condition for fitting expert data while the fake label $\mathbf{c}^{-}$ corresponds to agent data. We implement $\mathbf{c}^{+}$ as $\mathbf{1}$ and $\mathbf{c}^{-}$ as $\mathbf{0}$.
To approximate the expectation in Eq. 3, we use random sampling, allowing us to achieve the result with just a single denoising step. Subsequently, given a state-action pair $(\mathbf{s},\mathbf{a})$, $\mathcal{L}_{\text{diff}}(\mathbf{s},\mathbf{a},\mathbf{c}^{+})$ measures how well $(\mathbf{s},\mathbf{a})$ fits the expert distribution and $\mathcal{L}_{\text{diff}}(\mathbf{s},\mathbf{a},\mathbf{c}^{-})$ measures how well $(\mathbf{s},\mathbf{a})$ fits the agent distribution [^1] For simplicity, we will use the notations $\mathcal{L}^{+}_{\text{diff}}$ and $\mathcal{L}^{-}_{\text{diff}}$ to represent $\mathcal{L}_{\text{diff}}(\mathbf{s},\mathbf{a},\mathbf{c}^{+})$ and $\mathcal{L}_{\text{diff}}(\mathbf{s},\mathbf{a},\mathbf{c}^{-})$, respectively, in the rest of the paper. Additionally, $\mathcal{L}^{+,-}_{\text{diff}}$ denotes $\mathcal{L}^{+}_{\text{diff}}$ and $\mathcal{L}^{-}_{\text{diff}}$ given a state-action pair.. That said, given state-action pairs sampled from expert demonstration, $\mathcal{L}^{+}_{\text{diff}}$ should be close to 00, and $\mathcal{L}^{-}_{\text{diff}}$ should be a large value; on the contrary, given agent state-action pairs, $\mathcal{L}^{+}_{\text{diff}}$ should be a large value and $\mathcal{L}^{-}_{\text{diff}}$ should close to 00.
@@ 4.3 Diffusion-Reward Adversarial Imitation Learning
Among the related works, DiffAIL [59] is the closest to ours, as it also uses a diffusion model for adversarial imitation learning. DiffAIL employs an unconditional diffusion model to denoise state-action pairs from both experts and agents. However, this approach only implicitly reflects the likelihood of state-action pairs belonging to the expert class through diffusion loss, making it challenging to explicitly distinguish between expert and agent behaviors.
In contrast, our method, DRAIL, uses a conditional diffusion model that directly conditions real ($c^{+}$) and fake ($c^{-}$) labels. This allows our model to explicitly calculate and compare the probabilities of state-action pairs belonging to either the expert or agent class. This clearer and more robust signal for binary classification aligns more closely with the objectives of the GAIL framework, leading to more stable and effective learning. For further details and the mathematical formulation, please refer to Section A.
@@ 5.2 Baselines
  • Generative Adversarial Imitation Learning with Gradient Penalty (GAIL-GP) is an extension of GAIL that introduces a gradient penalty to achieve smoother rewards and stabilize the discriminator.
@@
  • Diffusion Adversarial Imitation Learning (DiffAIL) [58] integrates a diffusion model into AIL by using the diffusion model loss to provide reward $e^{-\mathcal{L}_{\text{diff}}}$.
@@ 5.3 Experimental results
We present the success rates (Maze, FetchPush, HandRotate, AntReach) and return (Walker, CarRacing) of all the methods with regards to environment steps in Figure 4. Each task was trained using five different random seeds. Note that BC and Diffusion Policy are offline imitation learning algorithms, meaning they cannot interact with the environment, so their performances are represented as horizontal lines. Detailed information on model architectures, training, and evaluation can be found in Section F and Section G.
@@
We empirically found that our proposed DRAIL is robust to hyperparameters, especially compared to GAIL and WAIL, as shown in the hyperparameter sensitivity experiment inSection D.
@@ 5.4 Generalizability
The results of FetchPush under 1×1$\times$1 ×, 1.25×1.25$\times$1.25 ×, 1.75×1.75$\times$1.75 ×, and 2.0×2.0$\times$2.0 × noise level are presented in Figure 5. Across all noise levels, approaches utilizing the diffusion model generally exhibit better performance. Notably, our proposed DRAIL demonstrates the highest robustness towards noisy environments. Even at the highest noise level of 2.00×2.$0$, DRAIL maintains a success rate of over $95\%$, surpassing the best-performing baseline, Diffusion Policy, which achieves only around a $79.20\%$ success rate. In contrast, DiffAIL experiences failures in $2$ out of the $5$ seeds, resulting in a high standard deviation (mean: $40.90$, standard deviation: $47.59$), despite our extensive efforts on experimenting with various configurations and a wide range of hyperparameter values.
The extended generalization experiments results in Maze, FetchPush, HandRotate , and the new task FetchPick are presented in Section C. Overall, our method outperforms or performs competitively against the best-performing baseline, demonstrating its superior generalization ability.
@@ 5.5 Data efficiency
Figure 7: Reward function visualization. We present visualizations of the learned reward values by the discriminative classifier of GAIL and the diffusion discriminative classifier of our DRAIL. The target expert demonstration for imitation is depicted in (a), which is a discontinuous sine function. The reward distributions of GAIL and our DRAIL are illustrated in (b) and (c), respectively.
@@ 6 Conclusion
This work proposes a novel adversarial imitation learning framework that integrates a diffusion model into generative adversarial imitation learning (GAIL). Specifically, we propose a diffusion discriminative classifier that employs a diffusion model to construct an enhanced discriminator, yielding more robust and smoother rewards. Then, we design diffusion rewards based on the classifier’s output for policy learning. Extensive experiments in navigation, manipulation, locomotion, and game justify our proposed framework’s effectiveness, generalizability, and data efficiency. Future work could apply DRAIL to image-based robotic tasks in real-world or simulated environments and explore its potential in various domains outside robotics, such as autonomous driving, to assess its generalizability and adaptability. Additionally, exploring other divergences and distance metrics, such as the Wasserstein distance or f-divergences, could potentially further improve training stability.
Acknowledgement
This work was supported by NVIDIA Taiwan Research &\&& Development Center (TRDC) under the funding code 112HT911007. Shao-Hua Sun was supported by the Yushan Fellow Program by the Ministry of Education, Taiwan. Ping-Chun Hsieh is supported in part by the National Science and Technology Council (NSTC), Taiwan under Contract No. NSTC 113-2628-E-A49-026. We thank Tim Pearce and Ching-An Cheng for the fruitful discussions.
@@ Appendix A Relation to DiffAIL
Experimental setup. Wang et al. [58] evaluated DiffAIL and prior methods in locomotion tasks; in contrast, our work extensively compares our proposed framework DRAIL with various existing methods in various domains, including navigation (Maze and AntReach), locomotion (Walker and AntReach), robot arm manipulation (FetchPush and FetchPick), robot arm dexterous (HandRotate), and games (CarRacing). Additionally, we present experimental results on generalization to unseen states and goals on the goal-oriented tasks, and on varying amounts of expert data.
@@ C.1 Experiment settings
To show our approach’s better generalization capabilities, we extend the environment scenarios following the setting stated in Lee et al. [33]: (1) In Maze main experiment, the initial and goal states of the expert dataset only constitute $50\%$ of the potential initial and goal states. In the generalization experiment, we gather expert demonstrations from some lower and higher coverage: $25\%$, $75\%$, and $100\%$. (2) In FetchPick, FetchPush, and HandRotate main experiments, the demonstrations are collected in a lower noise level setting, 1×1$\times$1 ×. Yet, the agent is trained within an environment incorporating 1.5×1.5$\times$1.5 × noise, which is 1.5 times larger noise than the collected expert demonstration, applied to the starting and target block positions. In the generalization experiment, we train agents in different noise levels: 1×1$\times$1 ×, 1.25×1.25$\times$1.25 ×, 1.75×1.75$\times$1.75 ×, 2.0×2.0$\times$2.0 ×. (3) In AntReach main experiment, no random noise is added to the initial pose during policy learning. In AntReach generalization experiment, we train agents in different noise levels: 00 (default), $0.01$, $0.03$, $0.05$.
@@ C.2 Experiment results
Figure 8: Extended results of generalization experiments. Maze is evaluated with different coverages of state and goal locations in the expert demonstrations, while FetchPick, FetchPush, and HandRotate environments are evaluated in environments of different noise levels. The number indicates the amount of additional noise in agent learning compared to that in the expert demonstrations, with more noise requiring harder generalization. The noise level rises from left to right.
@@
Appendix D Hyperparameter Sensitivity Experiment
We empirically found that our proposed method, DRAIL, is robust to hyperparameters and easy to tune, especially compared to GAIL and WAIL. In this section, we present additional ablation experiments to examine how hyperparameter tuning affects the performance of DRAIL.
Like most AIL methods, the key hyperparameters of DRAIL are the learning rates of the policy and discriminator. the key hyperparameters for DRAIL are the learning rates of the policy and discriminator. We experimented with various learning rate values, including 5x, 2x, 1x, 0.5x, and 0.2x of the value used in the main results for the Maze environment. The results, presented in Figure 9, demonstrate that our method is robust to variations in hyperparameters.
(a) Discriminator Learning Rate
(b) Policy Learning Rate
Figure 9: Hyperparameter Sensitivity of DRAIL in the Maze Environment. The results show the performance of DRAIL under varying learning rates for the discriminator (9(a)) and policy (9(b)). Different scaling factors (5x, 2x, 1x, 0.5x, 0.2x) of the baseline learning rate are tested. The results demonstrate that DRAIL remains robust across these variations, maintaining stable performance in the Maze environment.
Appendix E Converged performance